OpenStudy (anonymous):

Solve each system of equations using substitution. x – y = 1 x = 1/2y + 2 Select one: (2, 2) (3, 0) (3, 2) (0, 2)

OpenStudy (amoodarya):

put x=1/2y+2 in first equation simplify you will have value of y so find x

OpenStudy (anonymous):

ok

OpenStudy (amoodarya):

x = 1/2y + 2 x – y = 1 ----> (1/2y + 2)-y=1 multiply by 2

OpenStudy (amoodarya):

lets solve it together

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

Ok

OpenStudy (amoodarya):

y=?

OpenStudy (anonymous):

Y = idk :( thats really confusing

OpenStudy (amoodarya):

ok again put x = 1/2y + 2 into x – y = 1 so you have (1/2y+2) – y = 1 ok ?

OpenStudy (anonymous):

ok

OpenStudy (amoodarya):

now solve for y (1/2y+2) – y = 1 1/2y+2-y=1 1/2 y-y=1-2 1/2y-2/2y=-1 ok ?

OpenStudy (anonymous):

ok

OpenStudy (amoodarya):

-1/2 y=-1 so multiply by _2 y=+2 ok ?

OpenStudy (anonymous):

ok

OpenStudy (amoodarya):

now we find x before x = (1/2y+2) so (y=2) x=1/2(2)+2 x=1+2 x=3 ok ?

OpenStudy (anonymous):

ok so (3,2)

OpenStudy (anonymous):

Yay! Can i show u one more

OpenStudy (amoodarya):

yes

OpenStudy (anonymous):

y = x – 2 y = 1/2x + 4 Select one: (12, 5) (3, 5) (12, 10) (6, 10)

OpenStudy (amoodarya):

put y = x – 2 into second equation and write here please

OpenStudy (anonymous):

y = x-2

OpenStudy (anonymous):

is it 12, 10

OpenStudy (amoodarya):

from 1st we have y = x – 2 x-2=y = 1/2x + 4 *2 to be easier 2x-4=x+8 2x-x=4+8 x=12 you are right

OpenStudy (anonymous):

Yay thanks :)

OpenStudy (amoodarya):

your welcome